Session: Feb 2001 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Thu Mar 1 18:08:50 MET 2001) Minor failures
Problems at the start of the schedule. Lost the first hour.

Shanghai (Huang Xinyong at Fri Mar 2 01:47:54 MET 2001) Success

\(Other\) (A Polatidis at Fri Mar 2 08:59:14 MET 2001) Minor failures
The observation was made with a new HEMPT that gives a receiver temperature around 50K. Due to a faulty power supply, we probably have lost all observations after 1230UT and maybe a few chunks earlier. Apologies for this.

Metsahovi (Prisse Könönen at Fri Mar 2 09:47:11 MET 2001) Failure
We did not participate this session. Most important reasons were that we had no thin tapes for the session (having not received supplements) and that our personnel (few in number) is busy with testing a new receiver. Weather was snowy, too.

Jodrell_Bank (Alastair Gunn at Fri Mar 2 10:26:31 MET 2001) Success
No known problems.

Effelsberg (A. Kraus at Fri Mar 2 12:43:42 MET 2001) Minor failures
Because of snowfall we had to stop several times: UT0515-0600, 0730-0803, 0850-0911, 1405-1430.